# Recurrent neural network wave functions for Rydberg atom arrays on kagome lattice

### v1

@@ Abstract
Rydberg atom array experiments have demonstrated the ability to act as powerful quantum simulators, preparing strongly-correlated phases of matter which are challenging to study for conventional computer simulations. A key direction has been the implementation of interactions on frustrated geometries, in an effort to prepare exotic many-body states such as spin liquids and glasses. In this paper, we apply two-dimensional recurrent neural network (RNN) wave functions to study the ground states of Rydberg atom arrays on the kagome lattice. We implement an annealing scheme to find the RNN variational parameters in regions of the phase diagram where exotic phases may occur, corresponding to rough optimization landscapes. For Rydberg atom array Hamiltonians studied previously on the kagome lattice, our RNN ground states show no evidence of exotic spin liquid or emergent glassy behavior. In the latter case, we argue that the presence of a non-zero Edwards-Anderson order parameter is an artifact of the long autocorrelations times experienced with quantum Monte Carlo simulations. This result emphasizes the utility of autoregressive models, such as RNNs, to explore Rydberg atom array physics on frustrated lattices and beyond.
@@ II.1 Two dimensional RNNs
Figure 1: (a) An illustration of a positive RNN wave function. Each RNN cell receives an input $\bm{\sigma}_{n-1}$ and a hidden state $\bm{h}_{n-1}$ and outputs a new hidden state $\bm{h}_{n}$. This vector is taken as an input to the Softmax layer (denoted S) that computes the conditional probability $P_{i}$. (b) RNN autoregressive sampling scheme: after obtaining the probability vector $\bm{y}_{i}$ from the Softmax layer (S) in step $i$, we sample it to produce $\bm{\sigma}_{i}$. The latter is taken again as an input to the RNN along with the hidden state $\bm{h}_{i}$ to sample the following degree of freedom $\bm{\sigma}_{i+1}$. (c) Mapping of a Kagome lattice to a square lattice by embedding three atoms in a larger local Hilbert space. (d) A two-dimensional (2D) RNN with periodic boundary conditions for a $3\times 3$ lattice for illustration purposes. A bulk RNN cell receives two hidden states $\bm{h}_{i,j-1}$ and $\bm{h}_{i-(-1)^{j},j}$, as well as two input vectors $\bm{\sigma}_{i,j-1}$ and $\bm{\sigma}_{i-(-1)^{j},j}$ (not shown) as illustrated by the black solid arrows. RNN cells at the boundary receive additional hidden states $\bm{h}_{i,j+1}$ and $\bm{h}_{i+(-1)^{j},j}$, as well as two input vectors $\bm{\sigma}_{i,j+1}$ and $\bm{\sigma}_{i+(-1)^{j},j}$ (not shown), as demonstrated by the blue curved and solid arrows. The sampling path is taken as a zigzag path, as demonstrated by the dashed red arrows. The initial memory states of the 2D RNN and the initial inputs are null vectors, as indicated by the dashed black arrows.
@@
The atom configurations of a Rydberg atom array on a kagome lattice can be seen as an $L\times L\times 3$ array of binary degrees of freedom where $L$ is the size of each side of the lattice. As illustrated in Fig. 1(c), we can map our kagome lattice with a local Hilbert space of $2$ to a square lattice with an enlarged Hilbert space of size $82^{3}=$ 8 which we can study using our two-dimensional (2D) RNN wave function Hibat-Allah et al. (2023); Sprague and Czischek (2023).
@@
$\bm{h}_{i,j}$ is a memory state with two indices for each atom in the two-dimensional lattice. Here ‘Neighbours($\bm{\sigma}_{i,j}$)’ returns a concatenation of the neighbours of $\bm{\sigma}_{i,j}$. The same observation goes for ‘Neighbours($\bm{h}_{i,j}$)’. These neighbours correspond to incoming vectors indicated by the black and blue arrows as illustrated in Fig. 1(d). More specifically, we define
@@ II.2 Supplementing RNN optimization with annealing
The previous sum goes over all classical Rydberg configurations $\{\bm{\sigma}\}$ in the computational $z$-basis. Note that $S_{\rm classical}$ is a pseudo-entropy that can be efficiently estimated using our RNN wave function as opposed to the quantum von Neumann entropy. Additionally, $T(n)$ is a pseudo-temperature that is annealed from some initial value $T_{0}$ to zero as follows: $T(n)=T_{0}(1-n/N_{\rm a})$ where $n\in[0,N_{\rm a}]$ and $N_{\rm a}$ is the total number of annealing steps. We present more details about the hyperparameters of our training scheme in Appendix. C.
@@ II.3 Topological entanglement entropy
To investigate the existence of a topological property in the Rydberg atom arrays on the kagome lattice, we compute the topological entanglement entropy (TEE) Hamma et al. (2005a, b); Levin and Wen (2006); Kitaev and Preskill (2006); Flammia et al. (2009); Isakov et al. (2011); Wildeboer et al. (2017). For a gapped phase of matter, where the area law is satisfied, the Renyi-$2$ entanglement entropy follows the scaling law $S_{2}(A)=aL-\gamma+\mathcal{O}(L^{-1})$, assuming $A$ and $B$ is partition of the system, $L$ is the size of the boundary between $A$ and $B$ and $S_{2}(A)\equiv-\log(\text{Tr}(\rho^{2}_{A}))$. In this case, $\gamma$ is the so-called TEE. In this paper, we use the swap trick with our RNN wave function ansatz Hastings et al. (2010); Hibat-Allah et al. (2020); Wang and Davis (2020) to calculate the second Renyi entropy $S_{2}$ to extract the TEE $\gamma$.
@@ III Results
Figure 3: In these panels, we focus on the Blockade radius $R_{b}=1.95$. (a) Plot of two point correlations $\langle n_{\bm{0}}n_{\bm{r}}\rangle$ with $\delta=3.3$ and for a system size $N=8\times 8\times 3$ and $R_{c}=2$. (b) Plots of the topological entanglement entropy versus $\delta$ for two different values of the cutoff radius $R_{c}$, using the Levin-Wen (LW) construction and the Kitaev-Preskill (KP) construction, for $N=8\times 8\times 3$. (c) A plot of the Edwards-Anderson order parameter $q_{\text{EA}}$ defined in Eq. (6) as a function of $\delta$ for $N=8\times 8\times 3$. (d) A plot of the density overlaps $\frac{1}{N}\sum_{i=1}^{N}n_{i}^{(1)}n_{i}^{(2)}$ and the spin overlaps $\frac{1}{N}\sum_{i=1}^{N}S_{i}^{(1)}S_{i}^{(2)}$ at $\delta=3.3$. Here $S_{i}=2n_{i}-1$, and (1) and (2) are labels for two sets of samples obtained from our optimized RNN, that are aggregated from $10$ different training seeds, for $N=6\times 6\times 3$ and $R_{c}=2$. For each seed, we generate $2\times 10^{5}$ independent samples and divide them into two sets.
@@
where $N$ is the system size, $n_{i}$ is the occupation number of site $i$ and $\rho=(\sum_{i=1}^{N}n_{i})/N$. Deviations of this order parameter from zero values are signals of the existence of a spin-glass phase. In Fig. 3(c), we plot this order parameter as a function of $\delta$ with $R_{c}=2,4$ and $R_{b}=1.95$. We find that the values of the order parameter are consistent with zero, as opposed to the results of QMC in Ref. Yan et al. (2023). Furthermore, we report in Fig. 3(d) the density-density overlap $\frac{1}{N}\sum_{i=1}^{N}n_{i}^{(1)}n_{i}^{(2)}$ and the spin-spin overlap $\frac{1}{N}\sum_{i=1}^{N}S_{i}^{(1)}S_{i}^{(2)}$ between different RNN samples at $R_{b}=1.95,\delta=3.3$, and $R_{c}=2$. Here labels (1) and (2) correspond to two independent sets of samples, which are obtained from optimized RNNs with $10$ different training seeds. The Gaussian nature of the overlap distribution in both representations is another indicator that there is no static signature of a spin-glass order Castellani and Cavagna (2005).
The discrepancy in our results and previous QMC findings Yan et al. (2023) could be related to emergent glassy dynamics in the QMC simulations, which results in very long auto-correlations times and thus in a non-ergodic behavior. To corroborate our findings, we run QMC simulations Merali et al. (2023) for larger inverse temperatures compared to Ref. Yan et al. (2023), namely for $\beta\geq 200$ and using $2.2\times 10^{6}$ Monte Carlo samples. We find that the QMC prediction for the EA order parameter is given as $q^{\text{QMC}}_{\text{EA}}=0.0000018(5)$ for $R_{b}=1.95,\delta=3.3$, a system size $8\times 8\times 3$, and for a radius cutoff $R_{c}=2$. The previous result agrees very well with our RNN findings in Fig. 3(c). This result is also confirmed by the good agreement between the RNN energies and the QMC energies as shown in Appendix. D. Our findings are further supported by the results of Ref. Yan et al. (2022), which suggests the possibility of transition in a quantum dimer model between nematic to paramagnetic to staggered states. In conclusion, our numerical investigation suggests that the long auto-correlation time could be a limiting factor in the QMC results reported in Ref. Yan et al. (2023).
@@ IV Conclusions and Outlooks
Our main finding, suggested by the two-dimensional RNN wave functions results, points out that Rydberg atom arrays on the kagome lattice do not establish a $Z_{2}$ spin liquid in the highly entangled regime. This observation is also consistent with previous QMC studies Yan et al. (2023). Our RNN numerics also suggest that the highly entangled region corresponds to a trivial paramagnetic state and that there is no signature for spin glass order as opposed to the observations outlined in Ref. Yan et al. (2023). We believe that the ability of RNNs to generate uncorrelated samples from a multimodal distribution is a crucial factor in ruling out the spin-glass phase. Furthermore, supplementing RNNs with annealing turns out to be a valuable tool for mitigating local minima induced by the frustrated nature of the kagome lattice in the highly entangled regime. Additionally, we conclude that autocorrelation could be the main factor behind the observed spin glass phase observed in previous QMC simulations Yan et al. (2023).
Finally, we note that our method can be generalized to study other systems with potential topological order, such as the Rydberg atom arrays on the Ruby lattice Verresen et al. (2021); Semeghini et al. (2021); Giudici et al. (2022). One could also use quantum state tomography with RNNs Carrasquilla et al. (2019) in a wide variety of quantum simulators and also combine data with VMC to improve the variational results Bennewitz et al. (2021); Czischek et al. (2022); Moss et al. (2023). We also believe in the potential of RNN wave functions ansätzes in the discovery of new phases of matter with topological order. Overall, these results highlight the promising future of RNN wave functions Hibat-Allah et al. (2020); Roth (2020), language-model based wave functions, and neural quantum states Carleo and Troyer (2017) for investigating open questions and discovering new physics within the condensed matter community and beyond.
@@
Our code is made publicly available at “http://github.com/mhibatallah/RNNWavefunctions”. The hyperparameters we use are given in Appendix. C.
@@ Appendix C Hyperparameters
For all models studied in this paper, we note that for each annealing step, we perform $N_{\rm train}=5$ gradient steps. Concerning the learning rate $\eta$, we choose $\eta=10^{-3}$ during the warm-up phase and the annealing phase and switch to a learning rate $\eta=10^{-4}$ in the convergence phase. To train RNN on the $8\times 8\times 3$ lattices, we pre-train using the optimized RNN on the $6\times 6\times 3$ lattice without using annealing, since the RNN is expected to start from a variational energy that is close to the ground state energy in the new system size.
@@
Appendix D Numerical comparisons
In Tab. 2, we show a comparison between QMC’s and RNN’s energies per site for a system size $8\times 8\times 3$ and for a detuning $\delta=3.3$ and at the blockade radiuses $R_{b}=1.7,1.95$. These points correspond to the nematic and disordered phases, respectively. We note that our RNN-based ansatz provides energies with a relative error of less than $0.2\%$ compared to the QMC energies. The QMC simulations we run for Rydberg atom arrays are introduced in Ref. Merali et al. (2023). We use a finite-temperature QMC scheme run at several different values of $\beta$ until we observe convergence to the ground state. For each $\beta$, five independent simulations are taken and the convergence of observables is observed at $\beta=200$. Thus, to compute observables, we treat simulations with $\beta\geq 200$ as additional independent chains, giving us a total of 25 independent Markov chains at each parameter point. Each chain is allowed to warm-up for $10^{4}$ steps, after which $10^{6}$ sequential measurements were taken. With respect to the computation of the Edwards-Anderson order parameter, $q_{EA}$, we note that the analysis given in Ref. Yan et al. (2023) can give different results in the case of imperfect sampling. Ref. Yan et al. (2023) computes the order parameter independently for each Markov chain and then averages the results. This procedure can produce different results as each chain will only explore a subset of the QMC configuration space due to the presence of frustrated interactions. As a result, each chain’s estimate of the one-point function can be biased. Since $q_{EA}$ is a non-linear function of the one-point function, we must first aggregate the one-point functions generated by each Markov chain, and then compute $q_{EA}$. Lastly, to compute an estimate of the error in $q_{EA}$, we must account for auto-correlations and non-linearity simultaneously. This step is done by combining jackknife resampling with a binning procedure. To deal with auto-correlations, we first compute the one-point function on sequential “bins” of data; we found a bin size of $10^{4}$ to be sufficient, giving $100$ bins for each chain. Thus, we can consider each bin’s one-point function to be nearly uncorrelated, allowing us to directly apply the jackknife resampling procedure to these approximately independent bins.
To compare our RNN wave function ($d_{h}=100$) results with DMRG, we perform DMRG simulations using PastaQ Torlai and Fishman (2020) and ITensor Fishman et al. (2022) to further check the consistency of our RNN energies. In Tab. 3, we compare with DMRG using periodic boundary conditions. In the nematic phase $R_{b}=1.7,\delta=3.3$, we find an excellent match of the energies. For the disordered phase at $R_{b}=1.95,\delta=3.3$, our RNN energies are lower within an error of about $1\%$ and with orders of magnitude fewer parameters compared to DMRG. Furthermore, we choose the YC12 geometry used in Ref. Samajdar et al. (2021). We optimize our 2DRNN wave function ($d_{h}=60$) at $R_{b}=1.95$ and $\delta=3.5$. Our estimated energy is $-151.959(8)$ which is within 0.3% error compared to the DMRG energy provided in Ref. Samajdar et al. (2021).

### v2

@@ Abstract
Rydberg atom array experiments have demonstrated the ability to act as powerful quantum simulators, preparing strongly-correlated phases of matter which are challenging to study for conventional computer simulations. A key direction has been the implementation of interactions on frustrated geometries, in an effort to prepare exotic many-body states such as spin liquids and glasses. In this paper, we apply two-dimensional recurrent neural network (RNN) wave functions to study the ground states of Rydberg atom arrays on the kagome lattice. We implement an annealing scheme to find the RNN variational parameters in regions of the phase diagram where exotic phases may occur, corresponding to rough optimization landscapes. For Rydberg atom array Hamiltonians studied previously on the kagome lattice, our RNN ground states show no evidence of exotic spin liquid or emergent glassy behavior. In the latter case, we argue that the presence of a non-zero Edwards-Anderson order parameter is an artifact of the long autocorrelations times experienced with quantum Monte Carlo (QMC) simulations, and we show that autocorrelations can be systematically reduced by increasing numerical effort. This result emphasizes the utility of autoregressive models, such as RNNs, in conjunction with QMC, to explore Rydberg atom array physics on frustrated lattices and beyond.
@@ II.1 Two dimensional RNNs
Figure 1: (a) An illustration of a positive RNN wave function. Each RNN cell receives an input $\bm{\sigma}_{n-1}$ and a hidden state $\bm{h}_{n-1}$ and outputs a new hidden state $\bm{h}_{n}$. This vector is taken as an input to the Softmax layer (denoted S) that computes the conditional probability $P_{i}$. (b) RNN autoregressive sampling scheme: after obtaining the probability vector $\bm{y}_{i}$ from the Softmax layer (S) in step $i$, we sample it to produce $\bm{\sigma}_{i}$. The latter is taken again as an input to the RNN along with the hidden state $\bm{h}_{i}$ to sample the following degree of freedom $\bm{\sigma}_{i+1}$. (c) Mapping of a Kagome lattice to a square lattice by embedding three atoms in a larger local Hilbert space. (d) A two-dimensional (2D) RNN with periodic boundary conditions for a $\times 3$ 3 lattice for illustration purposes. A bulk RNN cell receives two hidden states $\bm{h}_{i,j-1}$ and $\bm{h}_{i-(-1)^{j},j}$, as well as two input vectors $\bm{\sigma}_{i,j-1}$ and $\bm{\sigma}_{i-(-1)^{j},j}$ (not shown) as illustrated by the black solid arrows. RNN cells at the boundary receive additional hidden states $\bm{h}_{i,j+1}$ and $\bm{h}_{i+(-1)^{j},j}$, as well as two input vectors $\bm{\sigma}_{i,j+1}$ and $\bm{\sigma}_{i+(-1)^{j},j}$ (not shown), as demonstrated by the blue curved and solid arrows. The sampling path is taken as a zigzag path, as demonstrated by the dashed red arrows. The initial memory states of the 2D RNN and the initial inputs are null vectors, as indicated by the dashed black arrows.
@@
The atom configurations of a Rydberg atom array on a kagome lattice can be seen as an $L\times L\times 3$ array of binary degrees of freedom where $L$ is the size of each side of the lattice. As illustrated in Fig. 1(c), we can map our kagome lattice with a local Hilbert space of 222 to a square lattice with an enlarged Hilbert space of size $2^{3}=8$ which we can study using our two-dimensional (2D) RNN wave function Hibat-Allah et al. (2023); Sprague and Czischek (2023).
@@
[…]
@@ II.2 Supplementing RNN optimization with annealing
The previous sum goes over all classical Rydberg configurations $\{\bm{\sigma}\}$ in the computational $z$-basis. Note that $S_{\rm classical}$ is a pseudo-entropy that can be efficiently estimated using our RNN wave function as opposed to the quantum von Neumann entropy. Additionally, $T(n)$ is a pseudo-temperature that is annealed from some initial value $T_{0}$ to zero as follows: $T(n)=T_{0}(1-n/N_{\rm a})$ where $n\in[0,N_{\rm a}]$ and $N_{\rm a}$ is the total number of annealing steps. We note that for each annealing step, we train our RNN for 555 training steps. We present more details about the hyperparameters of our training scheme in Appendix. C.
@@ II.3 Topological entanglement entropy
To investigate the existence of a topological property in the Rydberg atom arrays on the kagome lattice, we compute the topological entanglement entropy (TEE) Hamma et al. (2005a, b); Levin and Wen (2006); Kitaev and Preskill (2006); Flammia et al. (2009); Isakov et al. (2011); Wildeboer et al. (2017). For a gapped phase of matter, where the area law is satisfied, the Renyi-222 entanglement entropy follows the scaling law $S_{2}(A)=aL-\gamma+\mathcal{O}(L^{-1})$, assuming $A$ and $B$ is partition of the system, $L$ is the size of the boundary between $A$ and $B$ and $S_{2}(A)\equiv-\log(\text{Tr}(\rho^{2}_{A}))$. In this case, $\gamma$ is the so-called TEE. In this paper, we use the swap trick with our RNN wave function ansatz Hastings et al. (2010); Hibat-Allah et al. (2020); Wang and Davis (2020) to calculate the second Renyi entropy $S_{2}$ to extract the TEE $\gamma$.
@@ III Results
Figure 3: In these panels, we focus on the Blockade radius $R_{b}=1.95$. (a) Plot of two point correlations $\langle n_{\bm{0}}n_{\bm{r}}\rangle$ with $\delta=3.3$ and for a system size $N=8\times 8\times 3$ and $R_{c}=2$. (b) Plots of the topological entanglement entropy versus $\delta$ for two different values of the cutoff radius $R_{c}$, using the Levin-Wen (LW) construction and the Kitaev-Preskill (KP) construction, for $N=8\times 8\times 3$. (c) A histogram of the Edwards-Anderson order parameter $q_{\text{EA}}$ defined in Eq. (6) as a function of $\delta$ for $N=8\times 8\times 3$. (d) A plot of the density overlaps $\frac{1}{N}\sum_{i=1}^{N}n_{i}^{(1)}n_{i}^{(2)}$ and the spin overlaps $\frac{1}{N}\sum_{i=1}^{N}S_{i}^{(1)}S_{i}^{(2)}$ at $\delta=3.3$. Here $S_{i}=2n_{i}-1$, and (1) and (2) are labels for two sets of samples obtained from our optimized RNN, that are aggregated from 101010 different training seeds, for $N=6\times 6\times 3$ and $R_{c}=2$. For each seed, we generate $2\times 10^{5}$ independent samples and divide them into two sets. The error bars indicate the statistical uncertainty of one standard deviation, calculated across different samples. In this and other plots throughout the Article, error bars may be hidden if they are smaller than the symbol size.
@@
To address the finite-size scaling of the TEE, we compute the TEE at $R_{b}=1.95,\delta=3.3$ for $N=10\times 10\times 3$ by pre-training from the RNN optimized at $L=8$ and also for $L=12$ by starting from the optimized parameters at $L=10$. We follow a similar KP construction to Fig. 2 of the regions $A,B$ and $C$ where the size of each subregion is given as $\times 3\times 3$ 3, $\times 3\times 3$ 3, and 6×3×36$\times$ 3$\times$ 36 × 3 × 3 for $L=6$ respectively, and 4×4×34$\times$ 4$\times$ 34 × 4 × 3, 4×4×34$\times$ 4$\times$ 34 × 4 × 3, and 8×4×38$\times$ 4$\times$ 38 × 4 × 3 for $L=8$. Our estimates $\gamma_{\rm RNN}=0.10\pm 0.26,-0.1\pm 0.22$ for $L=6$ and $L=8$ respectively. These values, which are consistent with zero TEE, corroborate the absence of a $Z_{2}$ spin-liquid according to the RNN variational calculations. Note that the error bars can be systematically reduced by increasing the number of samples in the swap trick calculations Hibat-Allah et al. (2020, 2023).
@@
where $N$ is the system size, $n_{i}$ is the occupation number of site $i$ and $\rho=(\sum_{i=1}^{N}n_{i})/N$. Deviations of this order parameter from zero values are signals of the existence of a spin-glass phase. In Fig. 3(c), we plot this order parameter as a function of $\delta$ with $R_{c}=2,4$ and $R_{b}=1.95$. We find that the values of the order parameter are consistent with zero, as opposed to the results of QMC in Ref. Yan et al. (2023). Furthermore, we report in Fig. 3(d) the density-density overlap $\frac{1}{N}\sum_{i=1}^{N}n_{i}^{(1)}n_{i}^{(2)}$ and the spin-spin overlap $\frac{1}{N}\sum_{i=1}^{N}S_{i}^{(1)}S_{i}^{(2)}$ between different RNN samples at $R_{b}=1.95,\delta=3.3$, and $R_{c}=2$. Here labels (1) and (2) correspond to two independent sets of samples, which are obtained from optimized RNNs with 101010 different training seeds. The Gaussian nature of the overlap distribution in both representations is another indicator that there is no static signature of a spin-glass order Castellani and Cavagna (2005).
The discrepancy in our results and previous QMC findings Yan et al. (2023) could be related to emergent glassy dynamics in the QMC simulations, which results in very long auto-correlations times and thus in a non-ergodic behavior. To corroborate our findings, we run QMC simulations Merali et al. (2023), based on Stochastic Series Expansion (SSE) Sandvik and Kurkijärvi (1991); Sandvik (1999), for larger inverse temperatures compared to Ref. Yan et al. (2023), namely for $\beta\geq 200$ and using $2.2\times 10^{6}$ Monte Carlo samples. We find that the QMC prediction for the EA order parameter is given as $q^{\text{QMC}}_{\text{EA}}=0.0000018(5)$ for $R_{b}=1.95,\delta=3.3$, a system size 8×8×38$\times$ 8$\times$ 38 × 8 × 3, and for a radius cutoff $R_{c}=2$. The previous result agrees very well with our RNN findings in Fig. 3(c). This result is also confirmed by the good agreement between the RNN energies and the QMC energies as shown in Appendix. D. Our findings are further supported by the results of Ref. Yan et al. (2022), which suggests the possibility of transition in a quantum dimer model between nematic to paramagnetic to staggered states. In conclusion, our numerical investigation suggests that the long auto-correlation time could be a limiting factor in the QMC results reported in Ref. Yan et al. (2023).
@@
To indicate the quality of our variational calculations, we use the v-score as a metric Wu et al. (2024), which we report in Appendix F. Furthermore, to investigate the effect of parameter sharing in the RNN, illustrated in Eqs. (2), (3), we optimized our RNN using site-dependent parameters at $R_{b}=1.95,\delta=3.3$ and for $L=8$ and we find that both TEE and EA order parameters are consistent with zero within errors bars. This result confirms that parameter-sharing in our RNN ansatz does not bias our findings. More details are shared in Appendix G.
Finally, we report a comparison between RNN and QMC runtimes in Appendix H. In particular, we find a significant speed-up when training RNNs on an A100 GPU and QMC on a single CPU. This speedup is enabled by GPU hardware in addition to the ability of the 2DRNN to do transfer learning from smaller lattices to larger ones Roth (2020); Hibat-Allah et al. (2022, 2023); Luo et al. (2023); Moss et al. (2025a, b).
@@ IV Conclusions and Outlooks
Our main finding, suggested by the two-dimensional RNN wave functions results, points out that Rydberg atom arrays on the kagome lattice do not establish a $Z_{2}$ spin liquid in the highly entangled regime. This observation is also consistent with previous QMC studies Yan et al. (2023). Our RNN numerics also suggest that the highly entangled region corresponds to a trivial paramagnetic state and that there is no signature for spin glass order as opposed to the observations outlined in Ref. Yan et al. (2023). We believe that the ability of RNNs to generate uncorrelated samples from a multimodal distribution is a crucial factor for our numerics to indicate the non-existence of the spin-glass phase. Additionally, we find that autocorrelation could be the main factor behind the spin glass phase observed in previous QMC simulations Yan et al. (2023). In particular, our QMC numerics with more numerical effort compared to Ref. Yan et al. (2023) suggest the absence of a spin glass phase. Furthermore, supplementing RNNs with annealing turns out to be a valuable tool for mitigating local minima induced by the frustrated nature of the kagome lattice in the highly entangled regime. We highlight that advanced optimization schemes, such as minimal Stochastic Reconfiguration minSR Chen and Heyl (2024); Rende et al. (2024), is a potential avenue for enhancing the optimization of our 2DRNN ansätzes. Refs. Donatella et al. (2023); Lange et al. (2024) reported that Stochastic Reconfiguration (SR) is not as effective as Adam optimizer when applied to RNN wave functions. As a result, we believe that reconciling SR and RNN wave function optimization is an interesting research direction that deserves a thorough study in the future. Furthermore, studying dynamic properties of Rydberg atoms arrays is another promising research direction that can corroborate our findings by performing time-evolution on our RNN ansatz Sinibaldi et al. (2024); de Walle et al. (2024).
Finally, we note that our method can be generalized to study other systems with potential topological order, such as the Rydberg atom arrays on the Ruby lattice Verresen et al. (2021); Semeghini et al. (2021); Giudici et al. (2022). One could also use quantum state tomography with RNNs Carrasquilla et al. (2019) in a wide variety of quantum simulators and also combine data from QMC or quantum simulators with VMC to improve the variational results Bennewitz et al. (2021); Czischek et al. (2022); Moss et al. (2023); Lange et al. (2025). We also believe in the potential of RNN wave functions ansätzes in the discovery of new phases of matter with topological order. Overall, these results highlight the promising future of RNN wave functions Hibat-Allah et al. (2020); Roth (2020), language-model based wave functions, and neural quantum states Carleo and Troyer (2017) for investigating open questions and discovering new physics within the condensed matter community and beyond.
@@
Our code is made publicly available at “http://github.com/mhibatallah/RNNWavefunctions”. The hyperparameters we use are given in Appendix C.
Data availability
The data generated in this study is available from the corresponding author upon reasonable request.
@@ Appendix C Hyperparameters
For all models studied in this paper, we note that for each annealing step, we perform $N_{\rm train}=5$ gradient steps. Concerning the learning rate $\eta$, we choose $\eta=10^{-3}$ during the warm-up phase and the annealing phase and switch to a learning rate $\eta=10^{-4}$ in the convergence phase. To train RNN on the 8×8×38$\times$ 8$\times$ 38 × 8 × 3 lattices, we pre-train using the optimized RNN on the 6×6×36$\times$ 6$\times$ 36 × 6 × 3 lattice without using annealing, since the RNN is expected to start from a variational energy that is close to the ground state energy in the new system size.
@@
Appendix D Comparison with QMC and DMRG
In this appendix, we conduct a comparison between RNNs, QMC, and DMRG. First of all, we compare RNNs (with $N_{a}=4000$) and QMC from $R_{b}=0.5$ to 2.32.32.3 at fixed $\delta=3.3$ at a system size 8×8×38$\times$ 8$\times$ 38 × 8 × 3 and a cutoff radius $R_{c}=2$. We report our results in Fig. 5, where we find a very good agreement between our QMC and RNN energies per site $E/N$ and densities $\rho$. In particular, we observe that $\rho$ plateaus close to three different values, 1,1/31,1/31 , 1 / 3, and 1/61/61 / 6 respectively. The latter are consistent with previous findings Yan et al. (2023). In particular, $\rho\approx 1/3$ corresponds to the nematic phase and $\rho\approx 1/6$ to the highly-entangled disordered phase.
Figure 5: A comparison between RNN and QMC energies per site $E/N$ (left panel) and densities $\rho$ (right panel) for multiple values of $R_{b}$, fixed $\delta=3.3$ and $R_{c}=2$ on a system size 8×8×38$\times$ 8$\times$ 38 × 8 × 3.
[…]
Figure 6: Comparison between our RNN ansatz, QMC, and DMRG within the Nematic and paramagnetic (PM) phases and for different system sizes $L$ and cutoff radii $R_{c}$. Numerical values are reported in Tabs. 2,  3.
[…]